One year ago you invested $100,000 in a two-year asset that pays 4%, annualized (there is only 1 cash flow - after 2 years). a) what is the value of this asset today if the one-year (spot) interest rate is 4% and the two-year (spot) 7. rate is 5%? b) what is the value of this asset today if the one-year (spot) rate is 5% and the two-year rate is 4%?
Answer #1

The Equation used to solve the problem,

Value of the Asset (Today) = Cash Flow / Discounting Factor.

Return Cash Flow = (4% of $100,000) *2 = $8,000

Total Cash Flow = $8,000 + $100,000 = $108,000

a) Value Today = $ 108,000 / (1+0.04) = $103,846.1538

b) Value Today = $ 108,000 / (1+0.05) = $102,857.1429